A stock currently trading at $115 pays a $4 dividend in three months and nine months. An option on the stock with an exercise price of $105 expires in ten months. Annualized yield for T-bill for this option is 11% and annualized standard deviation (volatility) of the continuously compounded return on the stock is 17% per annum.     14. A stock currently trading at $115 pays a $4 dividend in three months and nine months. An option on the stock with an exer
Answer #1

a

Adjusted stock price = 115 – 4 * e^(-11%*3/12) - 4 * e^(-11%*9/12)

= 106.54

b , c, d

Option price= SN(d1) - Xe-r t N(d2)
d1 = [ ln(S/X) + ( r+ v2 /2) t ]/ v t0.5
d2 = d1 - v t0.5
Where
Current stock price= 115
Less: present value of dividend
Dividends 4 4
Paid in (months) 3 9
Present value of dividend                            -4.11        -4.34                -8.46
S= Stock price adjusted             106.54
X= Exercise price= 105
r= Risk free interest rate= 11.00%
v= Standard devriation= 17%
t= time to expiration (in years) =     0.8333
d1 = [ ln(106.544478845815/105) + ( 0.11 + (0.17^2)/2 ) *0.83333] / [0.17*0.83333^ 0.5 ]
d1 = [ 0.014602 + 0.103708 ] /0.155188
d1 =                                        0.762369
d2 = 0.762369 - 0.17 * 0.83333^0.5
                                       0.607181
N(d1) = N( 0.762369 ) =                      0.77708
N(d2) = N( 0.607181 ) =                      0.72813
Option price= 106.544478845815*0.777080018511513-105*(e^-0.11*0.83333) *0.728134491693128
                                              13.04

Price of call option is 13.04

Price of put option is:

Option price= = Xe –rt × N(-d2) – S × N(-d1)
d1 = [ ln(S/X) + ( r+ v2 /2) t ]/ v t0.5
d2 = d1 - v t0.5
Where
S= Current stock price=                         106.54
X= Exercise price= 105
r= Risk free interest rate= 11%
v= Standard devriation= 17%
t= time to expiration (in years)= 10/12 = 0.833333
d1 = [ ln(106.544478845815/105) + ( 0.11 + (0.17^2)/2 ) *0.83333] / [0.17*0.83333^ 0.5 ]
d1 = [ 0.0146 + 0.103708333333333 ] /0.155188
d1 = 0.7623687
d2 = 0.76237 - 0.17 * 0.83333^0.5
0.607180673
N(-d1) = N( - 0.76237 ) =                      0.22292
N(-d2) = N( - 0.60718 ) =                      0.27187
105 × e^(-0.11 × 0.83333) ×(1- N( 0.60718)) -106.544478845815× (1-N(0.76237))
Option price=                                      2.29

Price of put option is 2.29